# Long-Term Fairness Inquiries and Pursuits in Machine Learning: A Survey of Notions, Methods, and Challenges

### v2

@@ Abstract
The widespread integration of Machine Learning systems in daily life, particularly in high-stakes domains, has raised concerns about the fairness implications. While prior works have investigated static fairness measures, recent studies reveal that automated decision-making has long-term implications and that off-the-shelf fairness approaches may not serve the purpose of achieving long-term fairness. Additionally, the existence of feedback loops and the interaction between models and the environment introduces additional complexities that may deviate from the initial fairness goals. In this survey, we review existing literature on long-term fairness from different perspectives and present a taxonomy for long-term fairness studies. We highlight key challenges and consider future research directions, analyzing both current issues and potential further explorations.
@@ 1 Introduction
As Machine Learning (ML) algorithms assume increasingly influential roles in high-stake domains traditionally steered by human judgments, an extensive body of research has brought attention to the challenges of bias and discrimination against marginalized groups (Mehrabi et al., 2021; Cheng et al., 2021b). These issues are pervasive and manifest in different settings, including finance, legal (e.g., pretrial bail decisions), aviation, and healthcare practices, among others (Gohar et al., 2024; Barocas et al., 2023). The community has led tremendous research efforts toward measuring and mitigating algorithmic unfairness (Mehrabi et al., 2021). However, the majority of such fairness approaches quantify unfairness based on predefined statistical or causal criteria, often assuming a constant environment and system, defined as static fairness. This myopic conceptualization of fairness usually focuses on short-term outcomes and assesses only the instantaneous impact of interventions at a single snapshot, overlooking system dynamics (dynamic fairness) and/or long-term consequences over a time horizon (Zhang et al., 2020; Liu et al., 2018).
@@
Recent works have brought attention to the misalignment of long-term fairness considerations with the fairness objectives optimized in static settings and shown that imposing static fairness criteria often does not guarantee long-term fairness and may even amplify discrimination (Hu et al., 2019; Liu et al., 2018). One way static fairness approaches fall short is by failing to account for the future implications of current decisions on individuals or groups, undermining their effectiveness. For example, in predictive policing, Ensign et al. (2018) observed how an initial higher allocation of police in a specific area leads to more reported incidents, perpetuating increased surveillance and exacerbating the marginalization of those communities over time. To this end, a large body of work has proposed various methods to measure, mitigate, and evaluate different aspects of fairness in the long-term setting rather than achieving fairness for a single time step.
@@
Simply put, long-term fairness constitutes the settings outside of static fairness framework and short-term outcomes by addressing fairness over extended periods rather than focusing solely on immediate outcomes. While dynamic fairness aligns with this concept by considering evolving dynamics over time Li et al. (2023), long-term fairness has a much broader scope. This umbrella term has different facets, including sequential fairness (where sequential decisions impact fairness) and fairness over multiple time steps, among others (as depicted in Fig:1). In this work, we aim to unify the different strands of literature on long-term fairness under a common framework. Through this effort, we identify various dimensions and settings of long-term fairness and provide a rich review of existing research to explore diverse approaches and considerations for achieving fair outcomes in dynamic and non-static contexts.
@@
Difference from existing surveys. Several extensive surveys on ML fairness have been conducted. A subset of these surveys (Mehrabi et al., 2021; Gohar & Cheng, 2023; Pessach & Shmueli, 2022; Chouldechova & Roth, 2018; Finocchiaro et al., 2021) predominantly focus on static fairness or a specific vision of fairness (e.g., intersectional techniques), with only brief and cursory discussions on long-term fairness in some. Meanwhile, other works have incorporated some aspects of long-term fairness. Gajane et al. (2022) reviewed fairness approaches in Reinforcement Learning (RL), with overlap limited to very brief coverage of RL techniques for long-term fairness. Separately, Patro et al. (2022) critically examined the literature on Fair Ranking, discussing the limitations of static settings and exploring modeling frameworks, like simulations, within Ranking Systems only. The survey by Zhang & Liu (2021) focused only on sequential decision algorithms, representing one facet of long-term fairness inquiries, without a critical comparative analysis of definitions, methods, and frameworks. Finally, Tang et al. (2023b) reviewed the fairness concepts in ML, drawing connections to moral and political philosophy. They provide overviews of modeling frameworks and choices in dynamic settings without making them the primary focus. In comparison to these works, our survey specifically concentrates on long-term fairness and offers a comprehensive and up-to-date overview of the field, introducing the taxonomy encompassing long-term fairness notions, mitigation techniques, and evaluation methods. Additionally, we critically examine existing techniques, highlighting key challenges and open research questions. Our literature search methodology is detailed in Appendix A.
Contributions and Outline. Our main contributions are: (1) propose a taxonomy (Fig. 1) for long-term fairness, including notions, problem settings, mitigation approaches, and modeling frameworks, (2) examine long-term fair learning methods in multiple standard problem settings and evaluations, and (3) highlight key challenges and future directions. The rest of the paper is organized as follows. Section 2.3 introduces long-term fairness definition, problem settings, and modeling frameworks. Section 3 discusses fairness notions, while Section 4 covers mitigation techniques in dynamic settings. Section 5 outlines application-specific works, including Recommender Systems, and Section 6 explores alternative long-term fairness settings. Finally, Section 7 discusses open challenges and future directions.
@@ 2.1 What Is Long-Term Fairness?
Given the dynamic nature of real-world scenarios, long-term fairness emerges as a concept focused on achieving and maintaining fairness over a period of time rather than within isolated decisions or instances Hu & Zhang (2022). Long-term fairness has been described using various terms, often used interchangeably, including temporal, sequential, non-static, and dynamic fairness. However, these encompass different aspects of the overarching concept under complementary but orthogonal settings. Therefore, our first contribution is to identify and define these problem settings within a unified high-level taxonomy (Figure 1) for understanding the literature on long-term fairness.
@@ 2.2 Problem Setting
Long-term fairness was first studied within pipelines, where multiple decisions are required. For instance, a two-stage hiring model involves a candidate selection from the applicant pool, followed by hiring decisions Dwork & Ilvento (2018); Bower et al. (2017). However, long-term fairness has since expanded into a multifaceted concept, encompassing various problem settings. In Figure 1 (red box), we identify the four key problem settings that underpin inquiries into long-term fairness. Dynamic Modeling is the primary setting where the majority of the long-term fairness considerations have been explored. It addresses dynamic factors in the environment to maintain fairness over time, including cases of sequential fairness where fairness isn’t achieved over multiple decision rounds due to evolving system dynamics or the failure of static fairness notions when used at each time step Yin et al. (2023); Chouldechova & Roth (2018); Cheng et al. (2021a; 2022). Another facet that captures the real-world dynamics is the Distribution Shift setting, where deployment (testing) environments systematically differ from the training environment. This phenomenon can arise due to various factors, such as data non-stationarity (Barrainkua et al., 2023).
@@
Delayed impact is a special setting of long-term fairness where the impact of the decisions is not immediately observable but manifests gradually or after a significant time lag Liu et al. (2018). It focuses on a model of delayed outcomes, without which one cannot foresee the impact a fairness criterion would have if enforced as a constraint on a classification system. Finally, Performative Prediction refers to the causal influence that predictions have on their intended targets (Hardt & Mendler-Dünner, 2023). This phenomenon can trigger a self-reinforcing cycle where predictions shape future decisions, subsequently influencing outcomes. Perdomo et al. (2020) describe it as a form of distribution shift, but it uniquely focuses on distribution shift resulting solely from model deployment. While the impact of decisions on the environment has been explored in the dynamic modeling setting, performative prediction operates as a distinct framework (see Section 6). Apart from distribution shift, all these settings intersect with various techniques like bandits, reinforcement learning, performative prediction, and strategic classification, and hence, not clearly separable, making it challenging to write this survey and develop the taxonomy. To address this complexity and capture these interactions, we identify common modeling frameworks that accommodate these dynamics while offering flexibility in inquiry settings, as detailed below.
@@ 2.3 Dynamic Modeling For Long-Term Fairness
The study of long-term fairness dynamics in literature involves various approaches shaped by interactions with the environment and feedback mechanisms. Typically, these works (1) propose a dynamics model tailored to a specific domain, (2) expose unfairness from prolonged use of biased policies, and (3) suggest a “fair” policy to mitigate biases. As such, modeling frameworks (Fig: 1 green box) play a pivotal role in addressing the long-term fairness challenge. In this section, we outline commonly utilized frameworks for modeling dynamics in long-term fairness.
  • Pólya Urn: The Pólya Urn (Mahmoud, 2008) model simulates a stochastic process of drawing balls from an urn that increases the probability of observing certain events. At each time step, a ball is drawn and replaced with two of the same color, capturing the rich-get-richer dynamics Akpinar et al. (2022). It is a simple yet powerful framework to model self-reinforcing dynamics and cumulative advantages, making it particularly suitable for scenarios where preferential attachment dynamics Barabási & Albert (1999) are at play, e.g., resource allocation, social networks, and recommendation.
  • One-Step Feedback: It explores the impact of decision-making on the underlying population after only a single step of feedback, distinct from works that capture feedback over multiple time steps. It considers the dynamic response of the population at deployment, influencing fairness without model updates. This single-step evaluation is driven by practical considerations, acknowledging that other variables may dominate over extended periods in real-world scenarios Liu et al. (2018).
@@
  • Causal Modeling and Reasoning: Focusing on principled learning and reasoning about underlying data-generating processes, causality (Pearl, 2009; Guo et al., 2020) has been recognized as a unique tool for better understanding and mitigating bias in different contexts. Previous works have proposed to model and reason about causal relations among variables to characterize discrimination in static settings (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Chiappa & Isaac, 2018). Going beyond static towards dynamic and long-term settings, causal modeling of dynamics has also been explored in the literature to varying extents of detail (Creager et al., 2020; Zhang et al., 2020; Tang et al., 2023a).
Apart from these commonly used frameworks, relatively fewer works have also captured long-term fairness dynamics in social settings, (Heidari et al., 2019), online learning (Bechavod et al., 2019) strategic behavior Dong et al. (2018); Hu et al. (2019); Milli et al. (2019) and algorithmic recourse Ustun et al. (2019); Von Kügelgen et al. (2022).
@@ 3 Long-Term Fairness Notions
Static fairness notions often focus on immediate fairness, overlooking long-term equity risks, potentially exacerbating bias over time (D’Amour et al., 2020; Liu et al., 2018). In contrast, in the long-term fairness setting, the notions capture non-static, dynamic nuances of fairness over time. In this section, we present various long-term fairness notions used in literature in different settings, as shown in Figure 2. Some of these notions can still be used in static settings but have been adopted for long-term fairness using the dynamic modeling frameworks discussed in the previous section to capture dynamics. A summary of the main notations used in the paper is provided in the Appendix (Table 2).
Figure 2: Taxonomy of fairness notions used in long-term settings. The dotted boxes denote the adopted modeling frameworks
@@ 3.1 Effort-based Fairness
Static notions of fairness predominantly center on ensuring similar outcome probabilities across all protected groups. (Heidari et al., 2019; Mehrabi et al., 2021). However, these notions often operate under the assumption of a static dataset, treating individuals as immutable entities within the given data context. These approaches fail to consider that disadvantaged groups may need to exert more effort over time than advantaged groups to achieve similar outcomes, contradicting the principle of Equal Opportunity Roemer & Trannoy (2015); Von Kügelgen et al. (2022). This oversight persists because the distance to the decision boundary is typically greater for disadvantaged groups. Inspired by this, Heidari et al. (2019) proposed “effort unfairness”, which captures the disparity in average effort required by disadvantaged group members compared to the advantaged groups. It acknowledges that individuals can adjust their qualifications over time to achieve positive outcomes, highlighting its long-term and dynamic perspective on fairness. They propose Bounded-Effort unfairness as a measure to quantify this disparity while constraining it within a fixed level of effort. Given a data sample denoted by a tuple $(x,y)$ where $x\in\mathcal{X}$ $\subseteq\mathbb{R}^{d}$ and $y\in\mathcal{Y}=\{0,1\}$ denote the input vector and ground-truth label, respectively. Formally [^1] For ease of reading and comparison, we use notations defined in Guldogan et al. (2022) for all effort-based definitions.:
@@
Recognize that the condition $f(x)<0.5$ denotes an unfavorable outcome, signifying the individual’s unqualified status. In contrast, condition $f(x+\Delta x)\geq 0.5$ denotes a positive outcome, signifying successful effort $(\Delta x)$ to qualify the individual. For simplicity, the features are assumed to be numerical and monotone Duivesteijn & Feelders (2008). The parameter $\delta$ constraints the maximum effort to ensure that individuals or groups are not required to make disproportionately high efforts for improvement. In contrast to the Bounded-Effort approach, Threshold-Reward unfairness Heidari et al. (2019) shifts the focus from limiting effort to setting a lower bound on the reward. This perspective assesses how the minimum effort required to surpass a defined reward threshold varies among protected groups. Intuitively, this notion captures the idea that a disadvantaged group has to exert greater effort to get the same reward. Finally, Effort-Reward Heidari et al. (2019) unfairness is defined as the difference in maximum utility (difference between reward and effort) for each protected group after additional effort, avoiding reliance on specific thresholds to prevent fairness gerrymandering Kearns et al. (2018) and reduce result variability.
A closely related notion termed Equal Recourse quantifies the effort to attain desirable outcomes by measuring the average distance of individuals in different groups from the classifier’s decision boundary Gupta et al. (2019). Unlike bounded-effort unfairness, Equal Recourse aims to equalize average efforts across protected groups without imposing limits on effort.
@@
Finally, Huan et al. (2020) extends this framework to introduce a causal-based notion that enforces equality of effort. However, it does not capture the dynamics of the environment and model.

### v3

@@ Abstract
The widespread integration of Machine Learning systems in daily life, particularly in high-stakes domains, has raised concerns about the fairness implications. While prior works have investigated static fairness measures, recent studies reveal that automated decision-making has long-term implications and that off-the-shelf fairness approaches may not serve the purpose of achieving long-term fairness. Additionally, the existence of feedback loops and the interaction between models and the environment introduce additional complexities that may deviate from the initial fairness goals. In this survey, we review existing literature on long-term fairness from different perspectives and present a taxonomy for long-term fairness studies. We highlight key challenges and consider future research directions, analyzing both current issues and potential further explorations.
@@ 1 Introduction
As Machine Learning (ML) algorithms assume increasingly influential roles in high-stakes domains traditionally steered by human judgments, an extensive body of research has brought attention to the challenges of bias and discrimination against marginalized groups (Mehrabi et al., 2021; Cheng et al., 2021b). These issues are pervasive and manifest in different settings, including finance, legal (e.g., pretrial bail decisions), aviation, and healthcare practices, among others (Gohar et al., 2024; Barocas et al., 2023). The community has led tremendous research efforts toward measuring and mitigating algorithmic unfairness (Mehrabi et al., 2021). However, the majority of such fairness approaches quantify unfairness based on predefined statistical or causal criteria, often assuming a constant environment and system, defined as static fairness. This myopic conceptualization of fairness usually focuses on short-term outcomes and assesses only the instantaneous impact of interventions at a single snapshot, overlooking system dynamics (dynamic fairness) and/or long-term consequences over a time horizon (Liu et al., 2018; Zhang et al., 2020).
@@
Recent works have brought attention to the misalignment of long-term fairness considerations with the fairness objectives optimized in static settings, and have shown that imposing static fairness criteria often does not guarantee long-term fairness and may even amplify discrimination (Liu et al., 2018; Hu et al., 2019). One way static fairness approaches fall short is by failing to account for the future implications of current decisions on individuals or groups, undermining their effectiveness. For example, in predictive policing, Ensign et al. (2018) observed how an initial higher allocation of police in a specific area leads to more reported incidents, perpetuating increased surveillance and exacerbating the marginalization of those communities over time. To this end, a large body of work has proposed various methods to measure, mitigate, and achieve different aspects of fairness in the long-term setting rather than enforcing fairness for a single time step.
@@
Simply put, long-term fairness constitutes the settings outside of static fairness framework and short-term outcomes by addressing fairness over extended periods, rather than focusing solely on immediate outcomes. Importantly, it represents a spectrum of temporal fairness, from single-step effects (non-static) to sequential and multi-step fairness over extended periods, all of which are incorporated within our taxonomy (Figure 1) to capture the rich complexity of real-world scenarios. While dynamic fairness aligns with this concept by considering evolving dynamics over time (Li et al., 2023), long-term fairness has a much broader scope. This umbrella term has different facets, including sequential fairness (where sequential decisions impact fairness) and fairness over multiple time steps, among others (as depicted in Figure 1). In this work, we aim to unify the different strands of literature on long-term fairness under a common framework. Through this effort, we identify various dimensions and settings of long-term fairness and provide a rich review of existing research to explore diverse approaches and considerations for achieving fair outcomes in dynamic and non-static contexts.
@@
Difference from existing surveys.
Several extensive surveys on ML fairness have been conducted. A subset of these surveys (Chouldechova & Roth, 2018; Mehrabi et al., 2021; Finocchiaro et al., 2021; Pessach & Shmueli, 2022; Gohar & Cheng, 2023) predominantly focus on static fairness or a specific vision of fairness (e.g., intersectional techniques), with only brief and cursory discussions on long-term fairness in some. Meanwhile, other works have incorporated some aspects of long-term fairness. Gajane et al. (2022) reviewed fairness approaches in Reinforcement Learning (RL), with overlap limited to very brief coverage of RL techniques for long-term fairness. Separately, Patro et al. (2022) critically examined the literature on Fair Ranking, discussing the limitations of static settings and exploring modeling frameworks, like simulations, within Ranking Systems only. The survey by Zhang & Liu (2021) focused only on sequential decision algorithms, representing one facet of long-term fairness inquiries, without a critical comparative analysis of definitions, methods, and frameworks. Finally, Tang et al. (2023b) reviewed the fairness concepts in ML, drawing connections to moral and political philosophy. They provide overviews of modeling frameworks and choices in dynamic settings without making them the primary focus. In comparison to these works, our survey specifically concentrates on long-term fairness and offers a comprehensive and up-to-date overview of the field, introducing the taxonomy encompassing long-term fairness notions, mitigation techniques, and evaluation methods. Additionally, we critically examine existing techniques, highlighting key challenges and open research questions. Our literature search methodology is detailed in Appendix A.1.
Contributions and Outline.
Our main contributions are: (1) proposing a taxonomy (Figure 1) for long-term fairness, including notions, problem settings, mitigation approaches, and modeling frameworks, (2) examining long-term fair learning methods in multiple standard problem settings and evaluations, and (3) highlighting key challenges and future directions. The rest of the paper is organized as follows: Section 2 introduces long-term fairness definition, problem settings, and modeling frameworks. Section 3 discusses fairness notions, while Section 4 covers mitigation techniques in dynamic settings. Section 5 outlines application-specific works, and Section 6 explores alternative long-term fairness settings. Next, Section 7 presents a critical analysis of the literature, including methodological assumptions, and practical implications for evaluating long-term fairness.. Finally, Section 8 discusses open challenges and future directions.
@@ 2.1 What Is Long-Term Fairness?
Given the dynamic nature of real-world scenarios, long-term fairness emerges as a concept focused on achieving and maintaining fairness over a period of time rather than within isolated decisions or instances (Hu & Zhang, 2022). Long-term fairness has been described using various terms, often used interchangeably, including temporal, sequential, non-static, and dynamic fairness. However, these encompass different aspects of the overarching concept under complementary but orthogonal settings. Therefore, our first contribution is to identify and define these problem settings within a unified high-level taxonomy (Figure 1) for understanding the literature on long-term fairness. The taxonomy does not aim to enumerate or exhaust all possible combinations across the three dimensions identified, but rather serves as a flexible organizational scaffold for situating diverse approaches in the literature, with more specific interactions detailed in Figure 2 and Table 1.
@@ 2.2 Problem Setting
Long-term fairness was first studied within pipelines, where multiple decisions are required. For instance, a two-stage hiring model involves a candidate selection from the applicant pool, followed by hiring decisions (Dwork & Ilvento, 2018; Bower et al., 2017). However, long-term fairness has since expanded into a multifaceted concept, encompassing various problem settings. In Figure 1 (red box), we identify the four key problem settings that underpin inquiries into long-term fairness. Dynamic Modeling is the primary setting where the majority of the long-term fairness considerations have been explored. It addresses dynamic factors in the environment to maintain fairness over time, including cases of sequential fairness where fairness isn’t achieved over multiple decision rounds due to evolving system dynamics or the failure of static fairness notions when used at each time step (Chouldechova & Roth, 2018; Cheng et al., 2021a; 2022; Yin et al., 2023). It explicitly represents the temporal evolution of a system, where the current state depends on both past states and actions, as well as feedback from the environment or population. Another facet that captures the real-world dynamics is the Distribution Shift setting, where deployment (testing) environments systematically differ from the training environment. This phenomenon can arise due to various factors, such as data non-stationarity (Barrainkua et al., 2023).
@@
Delayed impact is a special setting of long-term fairness where the impact of the decisions is not immediately observable but manifests gradually or after a significant time lag (Liu et al., 2018). It focuses on a model of delayed outcomes, without which one cannot foresee the impact a fairness criterion would have if enforced as a constraint on a classification system. This framework is distinct because it emphasizes the lag between decisions and their eventual effects, typically without modeling ongoing or complex feedback from the environment. Finally, Performative Prediction is a recently introduced framework that formalizes scenarios involving the specific causal influence that predictions have on their intended targets (Hardt & Mendler-Dünner, 2023). Unlike dynamic modeling, which models how predictions influence future states over time, it abstracts away temporal dynamics and instead models the data distribution as a function of the deployed model itself. The goal is to find a performative equilibrium, a model that remains optimal even after influencing the data it observes. Perdomo et al. (2020) describe it as a form of distribution shift, but it uniquely focuses on distribution shift resulting solely from model deployment. While the impact of decisions on the environment has been explored in the dynamic modeling setting, performative prediction operates as a distinct framework (see Section 6). Apart from distribution shift, all these settings intersect with various techniques like bandits, reinforcement learning, performative prediction, and strategic classification, and hence, are not clearly separable, making it challenging to write this survey and develop the taxonomy. To address this complexity and capture these interactions, we identify common modeling frameworks that accommodate these dynamics while offering flexibility in inquiry settings, as detailed below.
@@ 2.3 Dynamic Modeling For Long-Term Fairness
The study of long-term fairness dynamics in literature involves various approaches shaped by interactions with the environment and feedback mechanisms. Typically, these works (1) propose a dynamics model tailored to a specific domain, (2) expose unfairness from prolonged use of biased policies, and (3) suggest a “fair” policy to mitigate biases. As such, modeling frameworks (Figure 1, green box) play a pivotal role in addressing the long-term fairness challenge. In this section, we outline commonly utilized frameworks for modeling dynamics in long-term fairness.
  • Pólya Urn: The Pólya Urn (Mahmoud, 2008) model is a uniquely powerful framework for capturing self-reinforcing dynamics. It simulates a stochastic process of drawing balls from an urn that increases the probability of observing certain events. At each time step, a ball is drawn and replaced with two of the same color, capturing the rich-get-richer dynamics (Akpinar et al., 2022). Its strength lies in modeling how small initial advantages can compound over time, while remaining mathematically simple, analytically tractable (Helfand, 2013), and highly interpretable. As a result, the Pólya urn is especially well-suited for studying cumulative advantage and preferential attachment in contexts like resource allocation, social networks, and recommendation systems (Barabási & Albert, 1999).
  • One-Step Feedback: It explores the impact of decision-making on the underlying population after only a single step of feedback, distinct from works that capture feedback over multiple time steps. It considers the dynamic response of the population at deployment, influencing fairness without model updates. This single-step evaluation is driven by practical considerations, acknowledging that other variables may dominate over extended periods in real-world scenarios (Liu et al., 2018).
@@
  • Causal Modeling and Reasoning: Focusing on principled learning and reasoning about underlying data-generating processes, causality (Spirtes et al., 1993; Pearl, 2009; Guo et al., 2020) has been recognized as a unique tool for better understanding and mitigating bias in different contexts. Previous works have proposed to model and reason about causal relations among variables to characterize discrimination in static settings (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Chiappa & Isaac, 2018). Going beyond static towards dynamic and long-term settings, causal modeling of dynamics has also been explored in the literature to varying extents of detail (Creager et al., 2020; Zhang et al., 2020; Tang et al., 2023a).
Other frameworks:
Apart from these commonly used frameworks, the literature has explored several less studied perspectives that capture different aspects of long-term fairness and its dynamics. A small but notable number of works examine incentive effects and strategic behavior (Hu et al., 2019), modeling how individuals may adapt or manipulate their features in anticipation of how an algorithm will respond, often using game-theoretic or best-response frameworks (e.g., Stackelberg games) (Dong et al., 2018; Milli et al., 2019; Zhang et al., 2022). Other less common frameworks include social settings (Heidari et al., 2019), online learning (Bechavod et al., 2019), and algorithmic recourse (Ustun et al., 2019; Von Kügelgen et al., 2022). While they are not yet major themes in the long-term fairness literature, they offer useful insights for understanding the broader dynamics and directions for future work.
@@ 3 Long-Term Fairness Notions
Static fairness notions often focus on immediate fairness, overlooking long-term equity risks, potentially exacerbating bias over time (D’Amour et al., 2020; Liu et al., 2018). In contrast, in the long-term fairness setting, the notions capture non-static, dynamic nuances of fairness over time. In this section, we present various long-term fairness notions used in literature in different settings, as shown in Figure 2. Some of these notions can still be used in static settings, but have been adopted for long-term fairness using the dynamic modeling frameworks discussed in the previous section to capture dynamics. A summary of the main notations used in the paper is provided in the Appendix (Table 2).
Figure 2: Taxonomy of fairness notions used in long-term settings. The dotted boxes denote the adopted modeling frameworks.
@@ 3.1 Effort-based Fairness
Static notions of fairness predominantly center on ensuring similar outcome probabilities across all protected groups (Heidari et al., 2019; Mehrabi et al., 2021). However, these notions often operate under the assumption of a static dataset, treating individuals as immutable entities within the given data context. These approaches fail to consider that disadvantaged groups may need to exert more effort over time than advantaged groups to achieve similar outcomes, contradicting the principle of Equal Opportunity (Roemer & Trannoy, 2015; Von Kügelgen et al., 2022). This oversight persists because the distance to the decision boundary is typically greater for disadvantaged groups. Inspired by this, Heidari et al. (2019) proposed “effort unfairness”, which captures the disparity in average effort required by disadvantaged group members compared to the advantaged groups. Crucially, effort-based fairness is inherently long-term and dynamic since it explicitly acknowledges that individuals can adjust their qualifications or features over multiple decision cycles, and that the effort required to reach a favorable outcome may accumulate over time. This perspective allows the assessment of fairness as an ongoing process, capturing disparities that persist or compound across repeated interactions with the system.
They propose Bounded-Effort unfairness as a measure to quantify this disparity while constraining it within a fixed level of effort. Given a data sample denoted by a tuple $(x,y)$ where $x\in\mathcal{X}$ $\subseteq\mathbb{R}^{d}$ and $y\in\mathcal{Y}=\{0,1\}$ denote the input vector and ground-truth label, respectively. Formally [^1] For ease of reading and comparison, we use notations defined in Guldogan et al. (2022) for all effort-based definitions.:
@@
Recognize that the condition $f(x)<0.5$ denotes an unfavorable outcome, signifying the individual’s unqualified status. In contrast, condition $f(x+\Delta x)\geq 0.5$ denotes a positive outcome, signifying successful effort $(\Delta x)$ to qualify the individual. For simplicity, the features are assumed to be numerical and monotone (Duivesteijn & Feelders, 2008). The parameter $\delta$ constraints the maximum effort to ensure that individuals or groups are not required to make disproportionately high efforts for improvement. In contrast to the Bounded-Effort approach, Threshold-Reward unfairness (Heidari et al., 2019) shifts the focus from limiting effort to setting a lower bound on the reward. This perspective assesses how the minimum effort required to surpass a defined reward threshold varies among protected groups. Intuitively, this notion captures the idea that a disadvantaged group has to exert greater effort to get the same reward. Finally, Effort-Reward (Heidari et al., 2019) unfairness is defined as the difference in maximum utility (difference between reward and effort) for each protected group after additional effort, avoiding reliance on specific thresholds to prevent fairness gerrymandering (Kearns et al., 2018) and reduce result variability.
A closely related notion termed Equal Recourse quantifies the effort to attain desirable outcomes by measuring the average distance of individuals in different groups from the classifier’s decision boundary (Gupta et al., 2019). Unlike bounded-effort unfairness, Equal Recourse aims to equalize average efforts across protected groups without imposing limits on effort.
@@
Finally, Huan et al. (2020) extends this framework to introduce a causal-based notion that enforces equality of effort. However, it does not capture the dynamics of the environment and the model.